OpenStudy (anonymous):

-3x+2y=8 what are possible answers for x and y

OpenStudy (anonymous):

x=-8/3 y=4

OpenStudy (lgbasallote):

many...infinite, in fact...unless you become more specific....

OpenStudy (lgbasallote):

^??

OpenStudy (anonymous):

lgbasallote is right

OpenStudy (anonymous):

what you described is a geometric line

OpenStudy (anonymous):

i think he wants the intercepts

OpenStudy (anonymous):

there's only one line he posted so it can't intercept

OpenStudy (anonymous):

\[y = 1/2 (3 x+8)\] is the best answer

OpenStudy (anonymous):

thank you all.